# Efficient Visual Transformer by Learnable Token Merging

### v1

@@ 3.2 Generating Token Merging Mask by Reducing Variational Upper Bound for the IB Loss
Given the training data $\left\{X_{i},y_{i}\right\}_{i=1}^{n}$ where $X_{i}$ is the $i$-the input training feature and $y_{i}$ is the corresponding class label. Let $Z_{i}$ be the the self-attention output tokens of the $X_{i}$, and $\tilde{X}_{i}(G)=\left(Z_{i}G\right)^{\top}$ is the merged tokens with $G$ being the token merging mask. We first specify how to compute the IB loss, $\textup{IB}(G)=I(\tilde{X}(G),X)-I(\tilde{X}(G),Y)$ which depends on $G$ and other network parameters, $X$ is a random variable representing the input feature which takes values in $\left\{X_{i}\right\}_{i=1}^{n}$, $\tilde{X}(G)$ is a random variable representing the merged tokens which takes values in $\left\{\tilde{X}_{i}(G)\right\}_{i=1}^{n}$. $Y$ is a random variable representing the class label which takes values in $\left\{y_{i}\right\}_{i=1}^{n}$. After performing K-means clustering on $\left\{\tilde{X}_{i}(G)\right\}_{i=1}^{n}$ and $\left\{X_{i}\right\}_{i=1}^{n}$, we have the clusters $\left\{\tilde{\cal C}_{a}\right\}_{a=1}^{C}$ and $\left\{{\cal C}_{b}\right\}_{b=1}^{C}$ for the merged tokens and the input features respectively, where $C$ is the number of classes. We also abbreviate $\tilde{X}(G)$ as $\tilde{X}$ for simplicity of the notations. Then we define the probability that $\tilde{X}$ belongs to cluster $\tilde{\cal C}_{a}$ as $\textup{Pr}\left[\tilde{X}\in a\right]=\frac{1}{n}\sum\limits_{i=1}^{n}\phi(\tilde{X}_{i},a)$ with $\phi(\tilde{X}_{i},a)=\frac{\exp\left(-{\left\|\tilde{X}_{i}-\tilde{\cal C}_{a}\right\|}_{2}^{2}\right)}{\sum_{a=1}^{A}\exp\left(-{\left\|\tilde{X}_{i}-\tilde{\cal C}_{a}\right\|}_{2}^{2}\right)}$. Similarly, we define the probability that $X_{i}$ belongs to cluster ${\cal C}_{b}$ as $\textup{Pr}\left[X\in b\right]=\frac{1}{n}\sum\limits_{i=1}^{n}\phi(X_{i},b)$. Moreover, we have the joint probabilities $\textup{Pr}\left[\tilde{X}\in a,X\in b\right]=\frac{1}{n}\sum\limits_{i=1}^{n}\phi(\tilde{X}_{i},a)\phi(X_{i},b)$ and $\textup{Pr}\left[\tilde{X}\in a,Y=y\right]=\frac{1}{n}\sum\limits_{i=1}^{n}\phi(\tilde{X}_{i},a){\rm 1}\kern-2.5pt{\rm I}_{\left\{y_{i}=y\right\}}$ where ${\rm 1}\kern-2.5pt{\rm I}_{\left\{\right\}}$ is an indicator function. As a result, we can compute the mutual information $I(\tilde{X}(G),X)$ and $I(\tilde{X}(G),Y)$ by $I(\tilde{X}(G),X)$ $=\sum\limits_{a=1}^{C}\sum\limits_{b=1}^{C}\textup{Pr}\left[\tilde{X}(G)\in a,X\in b\right]\log{\frac{\textup{Pr}\left[\tilde{X}(G)\in a,X\in b\right]}{\textup{Pr}\left[\tilde{X}(G)\in a\right]\textup{Pr}\left[X\in b\right]}},$ $I(\tilde{X}(G),Y)$ $=\sum\limits_{a=1}^{C}\sum\limits_{y=1}^{C}\textup{Pr}\left[\tilde{X}(G)\in a,Y=y\right]\log{\frac{\textup{Pr}\left[\tilde{X}\in a,Y=y\right]}{\textup{Pr}\left[\tilde{X}(G)\in a\right]\textup{Pr}\left[Y=y\right]}},$
and then compute the IB loss $\textup{IB}(G)$. As explained in our motivation, we aim to perform token merging while can reduce the IB loss. However, directly optimizing the IB loss in the standard SGD training is difficult as the IB loss is not separable. Given a variational distribution $Q(\tilde{X}\in a|Y=y)$ for $y,a\in[C]$ computed by Eq. (6.3) in Section 6.3, the following theorem gives a variational upper bound, $\textup{IBB}(G))$, for the IB loss $\textup{IB}(G))$. $\textup{IBB}(G))$ is separable and thus compatible with SGD training with minibatches.
@@
The proofs of Theorem 3.1 and Proposition 3.2 are deferred to Section 6. Inspired by Proposition 3.2, we can understand a transformer with token merging and multiple transformer blocks as an iterative process which reduces $\textup{IBB}(G)$ by gradient descent, where the $\ell$-th transformer block performs one-step gradient descent on $\textup{IBB}(G)$ according to (3.2). The mask module of at the $\ell$-th LTM-Transformer block generates the token merging mask $G^{(\ell)}$ from $G^{(\ell-1)}$, the token merging mask of the previous block, through (3.2). To improve the flexibility of the token merging mask, an MLP is applied on $Z_{i}$. Moreover, as IBB and $\nabla_{G}\textup{IBB}$ are separable, (3.2) can be performed on a minibatch ${\cal B}_{j}\subseteq\left\{1,\ldots,n\right\}$, which is compatible with minibatch-based training with SGD for a transformer network with LTM-Transformer blocks. That is, the mask module of the $\ell$-th LTM-Transformer block generates $G^{(\ell)}$ by
$G^{(\ell)}$ $=G^{(\ell-1)}$
$-\frac{2\eta}{n}\sum\limits_{i\in{\cal B}_{j}}\sum\limits_{a=1}^{C}\textup{MLP}(Z_{i})\frac{S_{i{a}}^{(l-1)}}{(\gamma_{i}^{(l-1)})^{2}}\left(\gamma_{i}^{(l-1)}{\cal C}_{a}-\zeta_{i}^{(l-1)}\right)\psi_{i,a},$ (3)
where $\textup{MLP}(\cdot)$ consists of two linear layers of the same hidden dimension with a ReLU activation in the middle, and each LTM-Transforer block has its own MLP.
@@
6:        Update $\phi(\tilde{X}_{i},a)$ for all the clusters $a\in[C]$ and $i\in[n]$.
@@
8:           Forward step: generate $\left\{G^{(\ell)}\right\}$ for all the LTM-Transformer blocks by (3.2) using the minibatch ${\cal B}_{j}$, the updated $\left\{\phi(\tilde{X}_{i},a)\right\}_{i\in{\cal B}_{j},a\in[C]}$, $\left\{Q^{(t-1)}(\tilde{X}\in a|Y=y)\right\}_{a\in[C],y\in[C]}$, and $\left\{\tilde{\cal C}^{(t-1)}_{a}\right\}_{a=1}^{C}$, as well as the output of the network
9:           Backward step: update the MLP layers of the mask modules in all the LTM-Transformer blocks as well as all the other weights in the neural network by a standard step of SGD on the cross-entropy loss
@@
11:        Compute $Q^{(t)}(\tilde{X}\in a|Y=y)$ by Eq. (6.3) in Section 6.3, perform K-means clustering on $\{\tilde{X}_{i}\}_{i=1}^{n}$ and update the clusters $\left\{\tilde{\cal C}^{(t)}_{a}\right\}_{a=1}^{C}$.
@@
Algorithm 1 describes the training process of a neural network with LTM-Transformer blocks using the standard cross-entropy loss for a classification problem. It is remarked that all the MLP layers of the mask modules in all the LTM-Transformer blocks, along with other network parameters, are updated with standard SGD. In order to generate the token merging masks for all the LTM-Transformer blocks before a new epoch starts, we update the variational distribution $Q^{(t)}$ and the clusters $\left\{\tilde{\cal C}^{(t)}_{a}\right\}_{a=1}^{C}$ at the end of the previous epoch.
@@ 4 Experimental Results
In this section, LTM-Transformers are assessed for the image classification task on the ImageNet-1k dataset. The results in Section 4.1 indicate that LTM outperforms existing state-of-the-art networks while maintaining a more compact architecture. In addition, LTM is compared with existing methods on token merging and shows better performance with lower computation costs. Furthermore, in Sections 4.2 and 4.3, we demonstrate that the use of LTM-MobileViT and LTM-EfficientViT as feature extraction backbones leads to superior mAP and reduced FLOPs compared to the baseline models for the tasks of object detection and semantic segmentation. In Section 4.4, we perform ablation studies on the effects of LTM-Transformer in reducing the IB loss and the IB bound at different layers of a LTM-Transformer network.
@@ 4.1 Image Classification
Implementation details. ImageNet [36] classification, we employ MobileViT-S [17], MobileViT-XS [17], EfficientViT-B1 [7], ViT-S [35], and Swin-T [4] as backbone architectures. We substitute the conventional transformer blocks in these backbones with LTM-Transformer blocks. During the training of LTM-Transformers, we utilize the AdamW optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The training process spans $300$ epochs, starting with a warm-up phase during which token merging is not applied in all the LTM-transformer blocks. After the warm-up stage, we enable token merging in all the LTM-transformer blocks. $t_{\text{warm}}$ is fixed to $100$ in all the experiments. All the experiments are conducted on four NVIDIA A100 GPUs with a total batch size of $1024$ images. Following prior works [4], our training incorporates popular data augmentation methods such as RandAugment, Mixup, Cutmix, and random erasing. We set the weight decay at $0.01$. The learning rate initially increases from $0.0002$ to $0.002$ over the first $10$ epochs and is subsequently reduced back to $0.0002$ following a cosine decay schedule. We set $\eta$ in (3.2) to $1$ in all the experiments. In addition, we apply a softmax operation on the token merging mask at each layer to ensure the aggregation weights for each merged tokens sum to $1$.
Figure 2: Top-1 accuracy vs FLOPs (G) on ImageNet-1k validation set.
@@
LTM-MobileViT-XS (Ours) 2.5 M 0.52 G 75.8
Mobile-Former 9.4 M 0.2 G 76.7
EfficientFormer [37] 12.3 M 1.3 G 79.2
@@
LTM-EfficientViT-B1 [r288] (Ours) 9.5 M 0.70 G 81.0
ViT-S/16 [3] 22.1 M 4.3 G 81.2
LTM-ViT-S/16 (Ours) 23.0 M 3.7 G 81.8
@@
LTM-Swin-T (Ours) 30.5 M 3.8 G 81.8
@@
TABLE I: Comparisons with baseline methods on ImageNet-1k validation set.
@@
Results. As shown in Table I, models integrated with LTM-Transformer blocks show reduced FLOPs and enhanced accuracy compared to their original visual transformer counterparts, at a cost of very slight increase of model size due to the extra parameters in the mask modules of the LTM-Transformer blocks. For instance, LTM-MobileViT-S not only reduces its FLOPs from $1.4$G to $1.17$G but also improves accuracy by $1.3\%$ over the original MobileViT-S, with a slight increase of the model size by $0.3$M. Similarly, LTM-MobileViT-XS achieves a $1.0\%$ accuracy increase while lowering its FLOPs from $0.7$G to $0.52$G compared to the original model, with a slight increase of model size by $0.2$M. To further demonstrate the efficiency of the LTM-Transformer, we compare it against current state-of-the-art weight pruning methods for efficient visual transformers, including S2ViTE [38], SPViT [39], and SAViT [40] on EfficientViT-B1 (r224). To apply S2ViTE, SPViT, and SAViT on EfficientViT-B1 (r224), we first run their pruning process following the standard implementation in their papers [38, 39, 40] on the ImageNet training data. After obtaining the pruned networks, we fine-tune them using the same setting as in [7]. As LTM-Transofmer focuses on token merging, we compare our LTM-Transformer models with existing token merging methods, including ToMe [14], ToFu [12], and LTMP [13]. All the token merging methods, ToMe [14], ToFu [12], LTMP [13], and our LTM-Transformer, are applied on the same visual transformer backbone. It is observed from Table I that with even lower FLOPs, LTM-Transformer models consistently outperform the competing token merging methods.
In addition, we apply LTM-Transformer to ViT-B with two different FLOPs budgets and compare LTM-Transformer with ViT-B models accelerated by ToMe, ToFu, and LTMP in Table II. The inference time of all the models are also evaluated on the validation set of ImageNet-1k and reported in milliseconds (ms) per batch for an evaluation batch size of 128 on one Nvidia A100 GPU. It is observed that LTM-Transformer models outperform the competing token merging methods with even lower FLOPs and faster inference speed. Remarkably, LTM-ViT-B with $12.85$G FLOPs has even higher top-1 accuracy than the vanilla ViT-B, evidencing the effectiveness of the informative token merging process in LTM-Transformer blocks.
Methods # Params. FLOPs Inference Time (ms/batch) Top-1
ViT-B 86.5 M 17.58 G 37.2 83.74
ViT-B (ToMe [14]) 86.5 M 8.78 G 25.4 78.88
ViT-B (ToFu [12]) 86.5 M 8.78 G 26.0 80.70
ViT-B (LTMP [13]) 86.5 M 8.84 G 27.1 81.21
LTM-ViT-B 90.0 M 8.30 G 25.2 82.23
ViT-B (ToMe [14]) 86.5 M 13.12 G 31.0 82.86
ViT-B (ToFu [12]) 86.5 M 13.12 G 31.5 83.22
ViT-B (LTMP [13]) 86.5 M 13.46 G 32.7 83.29
LTM-ViT-B 90.4 M 12.85 G 30.7 83.87
TABLE II: Performance comparison with token pruning methods on ImageNet.
4.2 Object Detection

### v2

@@ 3.2 Generating Token Merging Mask by Reducing the Variational Upper Bound for the IB Loss
Given the training data $\left\{X_{i},y_{i}\right\}_{i=1}^{n}$ where $X_{i}$ is the $i$-th input training feature and $y_{i}$ is the corresponding class label. Let $Z_{i}$ be the self-attention output tokens of the $X_{i}$, and $\tilde{X}_{i}(G)=\left(Z_{i}G\right)^{\top}$ is the merged tokens with $G$ being the token merging mask. We first specify how to compute the IB loss, $\textup{IB}(G)=I(\tilde{X}(G),X)-I(\tilde{X}(G),Y)$ which depends on $G$ and other network parameters, $X$ is a random variable representing the input feature which takes values in $\left\{X_{i}\right\}_{i=1}^{n}$, $\tilde{X}(G)$ is a random variable representing the merged tokens which takes values in $\left\{\tilde{X}_{i}(G)\right\}_{i=1}^{n}$. $Y$ is a random variable representing the class label which takes values in $\left\{y_{i}\right\}_{i=1}^{n}$. We first compute the class centroids $\left\{\tilde{\cal C}_{a}\right\}_{a=1}^{C}$ and $\left\{{\cal C}_{b}\right\}_{b=1}^{C}$ for the merged tokens and the input features by averaging the merged tokens and the input features in each class, where $C$ is the number of classes. We also abbreviate $\tilde{X}(G)$ as $\tilde{X}$ for simplicity of the notations. Then we define the probability that $\tilde{X}$ belongs to class $\tilde{\cal C}_{a}$ as $\textup{Pr}\left[\tilde{X}\in a\right]=\frac{1}{n}\sum\limits_{i=1}^{n}\phi(\tilde{X}_{i},a)$ with $\phi(\tilde{X}_{i},a)=\frac{\exp\left(-{\left\|\tilde{X}_{i}-\tilde{\cal C}_{a}\right\|}_{2}^{2}\right)}{\sum_{a=1}^{A}\exp\left(-{\left\|\tilde{X}_{i}-\tilde{\cal C}_{a}\right\|}_{2}^{2}\right)}$. Similarly, we define the probability that $X_{i}$ belongs to class ${\cal C}_{b}$ as $\textup{Pr}\left[X\in b\right]=\frac{1}{n}\sum\limits_{i=1}^{n}\phi(X_{i},b)$. Moreover, we have the joint probabilities $\textup{Pr}\left[\tilde{X}\in a,X\in b\right]=\frac{1}{n}\sum\limits_{i=1}^{n}\phi(\tilde{X}_{i},a)\phi(X_{i},b)$ and $\textup{Pr}\left[\tilde{X}\in a,Y=y\right]=\frac{1}{n}\sum\limits_{i=1}^{n}\phi(\tilde{X}_{i},a){\rm 1}\kern-2.5pt{\rm I}_{\left\{y_{i}=y\right\}}$ where ${\rm 1}\kern-2.5pt{\rm I}_{\left\{\right\}}$ is an indicator function. As a result, we can compute the mutual information $I(\tilde{X}(G),X)$ and $I(\tilde{X}(G),Y)$ by $I(\tilde{X}(G),X)$ $=\sum\limits_{a=1}^{C}\sum\limits_{b=1}^{C}\textup{Pr}\left[\tilde{X}(G)\in a,X\in b\right]\log{\frac{\textup{Pr}\left[\tilde{X}(G)\in a,X\in b\right]}{\textup{Pr}\left[\tilde{X}(G)\in a\right]\textup{Pr}\left[X\in b\right]}},$ $I(\tilde{X}(G),Y)$ $=\sum\limits_{a=1}^{C}\sum\limits_{y=1}^{C}\textup{Pr}\left[\tilde{X}(G)\in a,Y=y\right]\log{\frac{\textup{Pr}\left[\tilde{X}(G)\in a,Y=y\right]}{\textup{Pr}\left[\tilde{X}(G)\in a\right]\textup{Pr}\left[Y=y\right]}},$ and then compute the IB loss $\textup{IB}(G)$. As explained in our motivation, we aim to perform token merging while can reduce the IB loss. However, directly optimizing the IB loss in the standard SGD training is difficult as the IB loss is not separable. Given a variational distribution $Q(\tilde{X}\in a|Y=y)$ for $y,a\in[C]$ computed by Eq. (5) in Section 1.3 of the supplementary, the following theorem gives a variational upper bound, $\textup{IBB}(G)$, for the IB loss $\textup{IB}(G)$. $\textup{IBB}(G)$ is separable and thus compatible with SGD training with minibatches.
@@
The proofs of Theorem 3.1 and Proposition 3.2 are deferred to Section 1 of the supplementary. Inspired by Proposition 3.2, we can understand a transformer with token merging and multiple transformer blocks as an iterative process which reduces $\textup{IBU}(G)$ by gradient descent, where the $\ell$-th transformer block performs one-step gradient descent on $\textup{IBU}(G)$ according to (3.2). The mask module of at the $\ell$-th LTM block generates the token merging mask $G^{(\ell)}$ from $G^{(\ell-1)}$, the token merging mask of the previous block, through (3.2). To improve the flexibility of the token merging mask, an MLP is applied on $Z_{i}$. Moreover, as IBU and $\nabla_{G}\textup{IBU}$ are separable, (3.2) can be performed on a minibatch ${\cal B}_{j}\subseteq\left\{1,\ldots,n\right\}$, which is compatible with minibatch-based training with SGD for a transformer network with LTM blocks. In practice, the mask module of the $\ell$-th LTM block generates $G^{(\ell)}$ by
$\tilde{G}^{(\ell)}=$ $G^{(\ell-1)}$
$-\frac{2\eta}{n}\sum\limits_{i\in{\cal B}_{j}}\sum\limits_{a=1}^{C}Z_{i}\frac{S_{i{a}}^{(l-1)}}{(\gamma_{i}^{(l-1)})^{2}}\left(\gamma_{i}^{(l-1)}{\cal C}_{a}-\zeta_{i}^{(l-1)}\right)\psi_{i,a},$ (3)
$G^{(\ell)}=$ $\tilde{G}^{(\ell)}\circ M^{(\ell)}$ (4)
where $M^{(\ell)}\in[0,1]^{N\times P}$ is a binarized token merging mask generated by LTMP [14] for the $\ell$-th LTM block by applying the Gumbel-Softmax operation on $N\times P$ learnable parameters. The mask $M^{(\ell)}$ in our LTM serves as a learnable token merging mask module. Since the update formulation in Eq.(3) does not incorporate any trainable parameters, the number of the trainable parameters of an LTM block is the same as the number of the trainable parameters in a transformer block with LTMP, which is $N\times P$.
@@
6:        Update $\tau(\tilde{X}_{i},a)$ for all the classes $a\in[C]$ and $i\in[n]$.
@@
8:           Forward step: generate $\left\{G^{(\ell)}\right\}$ for all the LTM blocks by Eq. (3) using the minibatch ${\cal B}_{j}$, the updated $\left\{\tau(\tilde{X}_{i},a)\right\}_{i\in{\cal B}_{j},a\in[C]}$, $\left\{Q^{(t-1)}(\tilde{X}\in a|Y=y)\right\}_{a\in[C],y\in[C]}$, and $\left\{\tilde{\cal C}^{(t-1)}_{a}\right\}_{a=1}^{C}$, as well as the output of the network
9:           Backward step: update the MLP layers of the mask modules in all the LTM blocks as well as all the other weights in the neural network by a standard step of SGD on the cross-entropy loss
@@
11:        Compute $Q^{(t)}(\tilde{X}\in a|Y=y)$ by Eq. (5) in Section 1.3 of the supplementary, and update the class centroids $\left\{\tilde{\cal C}^{(t)}_{a}\right\}_{a=1}^{C}$.
@@
The training loss of our LTM models are the regular cross-entropy loss. The derived variational upper bound for the IB loss, IBB, is used to design the token mask module, where the token merging mask $G^{(\ell)}$ is generated by Eq. (3) and Eq. (4). Algorithm 1 describes the training process of a neural network with LTM-Transformer blocks using the standard cross-entropy loss for a classification problem. It is remarked that all the MLP layers of the mask modules in all the LTM-Transformer blocks, along with other network parameters, are updated with standard SGD. In order to generate the token merging masks for all the LTM-Transformer blocks before a new epoch starts, we update the variational distribution $Q^{(t)}$ at the end of the previous epoch.
Methods # Params. FLOPs Training Time Inference Time Top-1 Accuracy (%)
(minutes/epoch) (ms/batch) 0 1 5 10 25 50
MobileViT-XS [18] 2.3 M 0.70 G - 11.3 74.80 - - - - -
ToMe-MobileViT-XS [15] 2.3 M 0.54 G - 10.4 72.73 - - - - -
ToFu-MobileViT-XS [13] 2.3 M 0.54 G - 10.7 73.32 - - - - -
LTMP-MobileViT-XS[14] 2.3 M 0.56 G 16.0 10.9 - 73.91 73.79 73.98 74.05 74.18
LTM-MobileViT-XS (Fine-tuned) 2.3 M 0.52 G 16.5 10.7 - 74.25 74.31 74.54 74.70 74.95
MobileViT-S [18] 5.6 M 1.40 G - 15.1 78.40 - - - - -
ToMe-MobileViT-S [15] 5.6 M 1.22 G - 14.2 76.72 - - - - -
ToFu-MobileViT-S [13] 5.6 M 1.22 G - 14.4 77.24 - - - - -
LTMP-MobileViT-S[14] 5.6 M 1.26 G 18.1 14.5 - 77.53 77.69 77.82 78.03 78.14
LTM-MobileViT-S (Fine-tuned) 5.6 M 1.17 G 19.0 14.1 - 77.72 78.15 78.34 78.85 79.05
EfficientViT-B1 [r224] [6] 9.1 M 0.52 G - 10.0 79.40 - - - - -
ToMe-EfficientViT-B1 [r224] [15] 9.1 M 0.47 G - 9.6 78.81 - - - - -
ToFuEfficientViT-B1 [r224] [13] 9.1 M 0.47 G - 9.8 79.04 - - - - -
LTMP-EfficientViT-B1 [r224][14] 9.1 M 0.50 G 13.8 9.8 - 79.21 79.31 79.32 79.36 79.40
LTM-EfficientViT-B1 [r224] (Fine-tuned) 9.1 M 0.44 G 14.6 9.6 - 79.39 79.62 79.85 80.07 80.22
Swin-T [4] 29.0 M 4.50 G - 20.8 81.30 - - - - -
ToMe-Swin-T [15] 29.0 M 3.91 G - 17.5 79.28 - - - - -
ToFuSwin-T [13] 29.0 M 3.91 G - 17.8 79.65 - - - - -
LTMP-Swin-T [14] 29.0 M 3.95 G 21.0 17.9 - 79.78 79.96 80.09 80.24 80.30
LTM-Swin-T (Fine-tuned) 29.0 M 3.82 G 21.9 17.0 - 80.06 80.46 80.79 81.20 81.38
Swin-B [4] 88.0 M 15.4 G - 33.9 83.50 - - - - -
ToMe-Swin-B [15] 88.0 M 13.0 G - 29.9 81.87 - - - - -
ToFu-Swin-B [13] 88.0 M 13.0 G - 30.1 82.04 - - - - -
LTMP-Swin-B [14] 88.0 M 13.2 G 27.2 30.4 - 82.24 82.39 82.45 82.51 82.55
LTM-Swin-B (Fine-tuned) 88.0 M 12.0 G 28.9 29.6 - 82.50 82.72 82.88 83.43 83.64
ViT-S [35] 22.1 M 4.30 G - 22.5 81.20 - - - - -
ToMe-ViT-S [15] 22.1 M 3.82 G - 18.4 80.04 - - - - -
ToFu-ViT-S [13] 22.1 M 3.82 G - 18.7 80.19 - - - - -
LTMP-ViT-S [14] 22.1 M 3.89 G 21.4 19.0 - 80.32 80.40 80.35 80.41 80.50
LTM-ViT-S (Fine-tuned) 22.1 M 3.70 G 22.7 18.2 - 80.47 80.69 80.94 81.27 81.55
ViT-B [35] 86.5 M 17.58 G - 37.2 83.74 - - - - -
ToMe-ViT-B [15] 86.5 M 13.12 G - 31.0 82.86 - - - - -
ToFu-ViT-B [13] 86.5 M 13.12 G - 31.5 83.22 - - - - -
LTMP-ViT-B [14] 86.5 M 13.46 G 27.2 32.7 - 83.29 83.40 83.44 83.50 83.55
LTM-ViT-B (Fine-tuned) 86.5 M 12.85 G 28.4 30.7 - 83.35 83.57 83.76 83.91 83.96
TABLE I: Performance comparison between LTM with competing token merging baselines, ToMe [15], ToFu [13], and LTMP [14] in fine-tuning setup on ImageNet. Among the compared methods, ToMe and ToFu do not require training. Both LTM models and LTMP models are fine-tuned for 1, 5, 10, 25, and 50 epochs for fair comparisons.
@@ 4 Experimental Results
In this section, LTM-Transformers are assessed for the image classification task on the ImageNet-1k dataset. The results in Section 4.1 indicate that LTM outperforms existing state-of-the-art networks while maintaining a more compact architecture either fine-tuned from pre-trained backbones or trained from scratch. In addition, LTM is compared with existing methods on token merging and shows better performance with lower computation costs. Furthermore, in Sections 4.2 and 4.3, we demonstrate that the use of LTM-MobileViT and LTM-EfficientViT as feature extraction backbones leads to superior mAP and reduced FLOPs compared to the baseline models for the tasks of object detection and semantic segmentation. Comprehensive ablation studies are performed in Section 4.4. In Section 4.4.1, we study the impact of compression ratio on LTM. In Section 4.4.2, we perform ablation studies on the effects of LTM-Transformer in reducing the IB loss and the IB bound. In Section 4.4.3, we study IB loss and IB bound at different layers of an LTM-Transformer. In Section 4.4.4, we illustrate the training loss and the test loss in the training process of LTM-Transformers. In Section 4.4.5, we visualize merged tokens and their merging weights by LTM for selected images from ImageNet. In Section 4.4.6, we study the impact of compression ratio at different layers of the LTM-Transformer. Additional experiment results are presented in Section 2 of the supplementary of our paper. In Section 2.1 of the supplementary, we evaluate the transfer learning capability of the LTM-Transformer. In Section 2.2 of the supplementary, we evaluate the effectiveness of fine-tuning the LTM-Transformer with LoRA. In Section 2.3 of the supplementary, we evaluate the throughput of the LTM-Transformer with variant batch sizes.
@@ 4.1 Image Classification
Implementation details. In this section, we evaluate LTM models for ImageNet [36] classification. We employ MobileViT-S [18], MobileViT-XS [18], EfficientViT-B1 [6], ViT-S [35], ViT-B [35], Swin-T [4], and Swin-B [4] as backbone architectures. We substitute the conventional transformer blocks in these backbones with LTM blocks. All the experiments are conducted on four NVIDIA A100 GPUs with a total batch size of $1024$ images. Following prior works [4], our training incorporates popular data augmentation methods such as RandAugment, Mixup, Cutmix, and random erasing. We set $\eta$ in Eq. (3.2) to $1$ in all the experiments. In addition, we apply a softmax operation on the token merging mask at each layer to ensure the aggregation weights for each merged token sum to $1$. In all our experiments, we set the value of compression ratio $r=0.7$ for all our LTM models. We set the compression ratio of the compared methods to $0.75$ to see how LTM-Transformers compare with competing token merging methods with an even more significant reduction in the number of tokens. A study on the impact of the compression ratio $r$ on the performance of the LTM model is performed in Table V in Section 4.4.1.
We conduct the experiments of LTM for the token merging of vision transformers under two different training setups, which are the fine-tuning setup and the training-from-scratch setup. The experiments in the fine-tuning setup are conducted following the state-of-the-art token merging method, LTMP [14]. The training-from-scratch setup is designed to explore the potential of training LTM-Transformers from the beginning while reducing the IB loss with token merging, and the training with different backbones follows the same training settings as the original training process of the corresponding backbones [18, 6, 35, 4].
@@
LTM-MobileViT-XS (Ours) 2.3 M 0.52 G 75.8
@@
LTM-EfficientViT-B1 [r288] (Ours) 9.1 M 0.70 G 81.0
ViT-S [3] 22.1 M 4.3 G 81.2
LTM-ViT-S (Ours) 22.1 M 3.7 G 81.8
ViT-B [3] 86.5 M 17.6 G 83.7
LTM-ViT-B (Ours) 86.5 M 12.9 G 83.9
@@
LTM-Swin-T (Ours) 29.0 M 3.8 G 81.8
Swin-B [4] 88.0 M 15.4 G 83.5
LTM-Swin-B (Ours) 88.0 M 12.0 G 83.8
@@
TABLE II: Comparisons with baseline methods on ImageNet-1k validation set for the train-from-scratch setup.
Fine-Tuning Setup. Our proposed LTM can be straightforwardly applied to token merging with pre-trained models using the fine-tuning setup as in the existing state-of-the-art token merging method, LTMP [14]. In the fine-tuning setup, LTM models are not trained from scratch, and token merging for a pre-trained visual transformer is performed by simply changing all the transformer blocks of the pre-trained models to LTM-Transformer blocks. Following the settings in LTMP [14], the token merging mask modules are added to the original transformer blocks, and all the pre-trained weights are loaded as the initialization for the LTM models. In the fine-tuning process, the pre-trained weights are not updated and only the weights in the token merging mask modules, $\left\{M^{(\ell)}\right\}$, are updated. We fine-tune the LTM models for 1, 5, 10, 25, and 50 epochs, respectively, and compare them with LTMP models fine-tuned for the same number of epochs. Note that LTM models and LTMP models with the same backbones have the same number of parameters. It is observed that LTM models significantly outperform LTMP models fine-tuned for the same number of epochs. For example, the LTM-Swin-B fine-tuned for 50 epochs outperforms the LTMP-Swin-B fine-tuned for 50 epochs by $1.09\%$ in top-1 accuracy. The LTM-Swin-T fine-tuned for 50 epochs outperforms the LTMP-Swin-T fine-tuned for 50 epochs by $1.08\%$ in top-1 accuracy.
In addition, we also compare the LTM with three token merging methods, ToMe [15], ToFu [13], and LTMP [14], to demonstrate the superiority of our LTM. The results are shown in Table I. The inference time of all the models is also evaluated on the validation set of ImageNet-1k and reported in milliseconds (ms) per batch for an evaluation batch size of 128 on one Nvidia A100 GPU. We also evaluate the training cost of our LTM models compared with the other fine-tuning based token merging method, LTMP [14]. The training time is evaluated on 4 NVIDIA A100 GPUs with an effective batch size of 512 images. We report the average training time per epoch. The training overhead of LTM-Transformers mainly comes from the computation of $\left\{\phi(\tilde{X}_{i},a)\right\}_{i\in{\cal B}_{j},a\in[C]}$, $\left\{Q^{(t-1)}(\tilde{X}\in a|Y=y)\right\}_{a\in[C],y\in[C]}$, and $\left\{\tilde{\cal C}^{(t-1)}_{a}\right\}_{a=1}^{C}$ as described in Algorithm 1. It is observed that LTM only brings a marginal training time overhead compared to the competing token merging method.
@@
Training-from-Scratch Setup. As shown in Table II, LTM models show reduced FLOPs and enhanced accuracy compared to their original visual transformer counterparts. For instance, LTM-MobileViT-S not only reduces the FLOPs of MobileViT-S from $1.4$G to $1.17$G but also improves the top-1 accuracy by $1.3\%$. Similarly, LTM-Swin-T achieves a $0.5\%$ accuracy increase while lowering the FLOPs from $4.5$G to $3.7$G compared to the original Swin-T. Moreover, we compare LTM models against models compressed by current state-of-the-art weight pruning methods for efficient visual transformers, including S2ViTE [37], SPViT [27], and SAViT [29] on EfficientViT-B1 (r224). To apply S2ViTE, SPViT, and SAViT on EfficientViT-B1 (r224), we first run their pruning process following the standard implementation in their papers [37, 27, 29] on the ImageNet training data. After obtaining the pruned networks, we fine-tune the pruned models following the same setting as stated in their papers [37, 27, 29]. It is observed that LTM-EfficientViT-B1 outperforms all pruned models by at least $0.9\%$ in top-1 accuracy with even less FLOPs.